a) how many mL or carbon dioxide gas were generated by the decompostion of 6.24g of calcium carbonate at STP b) If 52.6 L of carbon dioxide at STP were needed, how many moles of calcium carbonate would be required?
Answer #1

CaCO3 --> CaO + CO2

PV = nRT

mol of CaCO3 = mass/MW = 6.24/100 = 0.0624 mol

1 mol = 22.4L alwas at STP

0.0624 mol --> 0.0624 *22.4 = 1.39776 L

b.

mol of CO2 = 52.6/22.4 = 2.3482

ratio is 1:1

so

2.3482 mol of CaCO3 needed